# Rethinking Model Prototyping through the MedMNIST+ Dataset Collection

### v2

@@ 4 Discussion and Conclusion
Table 15: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC) and area under the receiver operating characteristic curve (AUC) for the PneumoniaMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Table 16: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC) and area under the receiver operating characteristic curve (AUC) for the RetinaMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Table 17: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC) and area under the receiver operating characteristic curve (AUC) for the TissueMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.

### v3

@@ Acknowledgements
Supplementary Table S10: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC), for a fixed operating point of $0.5$ and area under the receiver operating characteristic curve (AUC) for the PneumoniaMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Supplementary Table S11: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC), for a fixed operating point of $0.5$ and area under the receiver operating characteristic curve (AUC) for the RetinaMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Supplementary Table S12: Benchmark outcomes summarizing the mean and standard deviation of accuracy (ACC), for a fixed operating point of $0.5$ and area under the receiver operating characteristic curve (AUC) for the TissueMNIST dataset across all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.